What is the radius of a right circular cone with a volume of 800 cubic inches and a height of 12 inches? Round your answer to the nearest hundredth...
ANSWER CHOICES ARE:
A) 7.98 inches
B) 8 inches
C) 7.98 square inches
D) 14.14 inches
Thanks for helping
-s.j.a.
: What is the radius of a right circular cone with a volume of 800 cubic inches and a height of 12 inches? Round your answer to the nearest hundredth...
ANSWER CHOICES ARE:
A) 7.98 inches
B) 8 inches
C) 7.98 square inches
D) 14.14 inches
Thanks for helping
-s.j.a.

Answer by checkley71(8405)

V=1/3PIR^2H
800=1/3*3.14*R^2*12
800=12.56*R^2
R^2=800/12.56
R^2=63.69
R=SQRT63.69
R=7.98 INCHES. A) IS THE CHOICE.
Question 117120: What is the radius of a right circular cone with a volume of 800 cubic inches and a height of 12 inches? Round your answer to the nearest hundredth...
ANSWER CHOICES ARE:
A) 7.98 inches
B) 8 inches
C) 7.98 square inches
D) 14.14 inches
Thanks for helping
-s.j.a.
: What is the radius of a right circular cone with a volume of 800 cubic inches and a height of 12 inches? Round your answer to the nearest hundredth...
ANSWER CHOICES ARE:
A) 7.98 inches
B) 8 inches
C) 7.98 square inches
D) 14.14 inches
Thanks for helping
-s.j.a.

Answer by near2u_28@yahoo.com(2)

1/3 pi r2h= 800
pi=22/7
r=radius
h=height
r2=200*7/22
r2=63 approx.
r=radius=7.98 cm